OpenStudy (anonymous):

Find the coordinates of the center and the length of the radius of the following circle. ( x - 3 ) 2 + y 2 = 4 center = (,) radius =

OpenStudy (ash2326):

you try this on your own, compare with the standard equation and find the center and radius

OpenStudy (anonymous):

i got (9,0) and 2 for radius

OpenStudy (ash2326):

see it's (x-a)^2+(y-b)^2=r^2 we have this ( x - 3 ) 2 + y 2 = 4 center will be 3,0 radius is 2